Cherokee Inc. is a merchandiser that provided the following information: Number of units sold Selling price per unit Variable selling expense per unit Variable administrative expense per unit Total fixed selling expense Total fixed administrative expense Beginning merchandise inventory Ending merchandise inventory Merchandise purchases 12,000 S 16.00 $ 2.0 S 1.00 $ 19,000 S 15,000 $ 11,000 S 24,000 $ 87,000 Required: 1. Prepare a traditional income statement. CHEROKEE, INC. Traditional Income Statement Selling and administrative expenses:
2. Prepare a contribution format income statement CHEROKEE, INC Contribution Format Income Statement Variable expenses: Fixed expenses:
Answer #1
1
Traditional Income Statement
Sales 192000 =12000*16
Cost of goods sold 74000 =11000+87000-24000
Gross margin 118000
Selling and administrative expenses:
Selling expenses 43000 =19000+(12000*2)
Administrative expenses 27000 =15000+(12000*1)
70000
Net operating income 48000
2
Contribution Format Income Statement
Sales 192000
Variable expenses:
Cost of goods sold 74000 =11000+87000-24000
Selling expenses 24000 =12000*2
Administrative expenses 12000 =12000*1
110000
Contribution margin 82000
Fixed expenses:
Selling expenses 19000
Administrative expenses 15000
34000
Net operating income 48000